OpenStudy (anonymous):

Task 2 Bernard is realizing that his dimensions are not fitting the way that he wants it to because he forgot about the deck. The deck is going to be a rectangular space that borders the pool. Also, for the pool to fit just right there must be a special product represented when either the area or volume expressions are determined. He wants you to design your own rectangular pool using polynomial expressions, including the deck to borders the pool.

OpenStudy (anonymous):

A. Include a diagram or sketch of your pool and deck. Your diagram must have all dimensions represented with polynomials and ensure that a special product will be used. B. Determine the polynomial expression to represent the area of land space that the pool will cover. C. Determine the area of the deck that borders the pool based on your dimensions. Be sure to include all of your work and calculations.

OpenStudy (perl):

no dimensions are given?

OpenStudy (anonymous):

x represents the width of the pool the length is 3*x+3=3xx+3 and the depth is 2*length -7 = 2(3x+3)-7 = 6x-1

OpenStudy (anonymous):

@perl

OpenStudy (anonymous):

need help?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

k

OpenStudy (anonymous):

which one are you looking for? Area or volume?

OpenStudy (anonymous):

pop this to the top

OpenStudy (anonymous):

of the list i mean

OpenStudy (anonymous):

GTG

OpenStudy (anonymous):

Tell me when your on again ima close it

OpenStudy (anonymous):

close to what?

OpenStudy (anonymous):

Closing the question